Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2018-11-16

Observers: M. Berton, E. Congiu
Support: A. Pastorello, A. Reguitti
TAT: R. Cartier, J. Chen

Standard EFOSC calibrations - biases, dome flats, and arcs (evening and morning)

Weather:
Start - thin clouds, seeing - 1.2
Middle - thin clouds, seeing - 1.2
End - clear, seeing ~1.3

Issues:
SOFI - usual problem with the sensor. Solved relatively quickly, but we lost enough time that we had to give up a target.
We repeated the SN2018hle OB to fill a hole in the schedule caused by the problem.

Instrument - EFOSC


UT            Target                      Type                Setup                  airmass             seeing              notes
---------------------------------------------------------------------------------------
24.20       Feige110                STD                all                        1.1                    1.0             
24.45       AT2018ids             class                gr13                    1.7                    1.4                    Ia
24.55       SN2018ilu             follow               gr13                    1.2                    1.2
01.25       AT2018ihx             class               gr13                     1.2                    1.2                    Ia
01.45       SN2017bzc           follow              BV                        1.1                    1.1                  
03.05       SN2018aad          follow               gr16                     1.3                    1.2

Instrument - SOFI
UT            Target                      Type                Setup                  airmass             seeing              notes
---------------------------------------------------------------------
04.05      Image analysis
04.30      SN2018aad            follow              Ks                       1.2                      0.9                    noise on the sensor, we managed to solve but we had to start the OB from the beginnning
05.20      SN2018hle             follow              JH                       1.1                      0.8                
05.50      GD71                       STD                BG                       1.5                     1.3
06.00      Telescope focus
06.10      SN2017ijn               follow              JHK                    2.4                      1.3
06.40      HIP48693                TELL               BG                     1.9                       1.3
06.42      SN2017ijn               follow              BG                      1.9                      1.3                   noise on the detector, OB started again
08.33     HIP48693                TELL               BG                     1.9                       1.3
   






Comments